OpenStudy (anonymous):

r=p-k ln t, solve for t can someone please explain how ln becomes e? thanks!

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

First isolate \(\ln t\)

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

when I enter r=p-klnt in my Mathway app, the answer comes with t=e^[(p-r)/k] but how does the ln t become e?

OpenStudy (anonymous):

Hey, can you solve for \(\ln t\) first? Then I can explain.

OpenStudy (anonymous):

Suppose you let \(\ln t = x\). Can you solve for \(x\)?\[ r=p-kx \]

OpenStudy (anonymous):

Yes or no? I'm running out of time.

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

So you can't even do basic algebra? Is that how I am supposed to interpret your silence?

OpenStudy (anonymous):

Okay all I can really tell you is that \[ b^a=c\iff \log_{b}(c)=a \]

OpenStudy (anonymous):

If we let \(b=e\), then \(\log_3(c) = \ln(c)\).\[ e^a=c\iff \ln(c)=a \]

OpenStudy (anonymous):

If you let \(c=t\), then we can see that \[ e^a=t\iff \ln(t) =a \]

OpenStudy (anonymous):

Finally, let \(a\) be whatever you got when you solved for \(\ln(t)\).

OpenStudy (anonymous):

sorry for not replying quicker...thanks for your help!